# ConDiSim: Conditional Diffusion Models for Simulation-Based Inference

### v1

@@ A.11 Hodgkin-Huxley Model
[…]
A.12 High-Dimensional Genetic Oscillator
The genetic oscillator model describes a biochemical reaction network involving $9$ species governed by $18$ reactions, parameterized by $15$ reaction rate constants. This network is based on a positive–negative feedback loop, where the activator protein $A$ upregulates transcription of both its gene and the repressor protein $R$. The repressor $R$ sequesters $A$ by forming a complex $C$, thereby inhibiting further activation [Vilar et al., 2002]. The species in the network are:
@@ A.12 High-Dimensional Genetic Oscillator
The network includes the following key reactions:
$D_{A}^{*}$ $\xrightarrow{\theta_{A}}D_{A},$ $D_{R}^{*}$ $\xrightarrow{\theta_{R}}D_{R},$
$D_{A},A$ $\xrightarrow{\gamma_{A}}D_{A}^{*},$ $D_{R},A$ $\xrightarrow{\gamma_{R}}D_{R}^{*},$
$D_{A}^{*}$ $\xrightarrow{\alpha_{a}^{\prime}}M_{A},$ $D_{R}^{*}$ $\xrightarrow{\alpha_{r}^{\prime}}M_{R},$
$M_{A}$ $\xrightarrow{\delta_{ma}}\phi,$ $M_{R}$ $\xrightarrow{\delta_{mr}}\phi,$
@@
    $\alpha_{a}$ $=50,\quad\alpha_{a}^{\prime}=500,\quad\alpha_{r}=0.01,\quad\alpha_{r}^{\prime}=50,$
    $\beta_{a}$ $=50,\quad\beta_{r}=5,\quad\delta_{ma}=10,\quad\delta_{mr}=0.5,$
    $\delta_{a}$ $=1,\quad\delta_{r}=0.2,\quad\gamma_{a}=1,\quad\gamma_{r}=1,$
    $\gamma_{c}$ $=2,\quad\theta_{a}=50,\quad\theta_{r}=100.$
@@
[…]
Appendix B Hyperparameters
In this section, we describe the hyperparameters and simulation environments employed across all methods. Training data for the SBI benchmark problems were generated using the sbibm [Lueckmann et al., 2021], except for the LV and SIR models, which were generated with the BayesFlow framework [Radev et al., 2023b]. For the HH model, we adopt the implementation provided by [Gloeckler et al., 2024], and for the Genetic Oscillator model, we follow the approach of [Akesson et al., 2022]. In this model, the simulated data is high dimensional, with a shape of $(N\times 3\times 200)$ representing three species observed over 200 timesteps. We compress this data using an autoencoder whose latent space is set to match the number of simulator parameters. For Simformer, we use hyperparameters published in their original works [Gloeckler et al., 2024]. For GATSBI, task-specific hyperparameters were taken from [Ramesh et al., 2022] and adjusted as needed based on each problem’s structure and computational demands.
@@ Appendix B Hyperparameters
For NPE, we follow [Lueckmann et al., 2021], which evaluated NPE across ten benchmark problems. As reported in Appendix H of sbibm paper, Neural Spline Flows (NSF) generally outperform Masked Autoregressive Flows (MAF), originally recommended for NPE. Consequently, we use the NSF implementation from the sbi package [Tejero-Cantero et al., 2020] as the density estimator in all experiments. The implementation of Automatic Posterior Transformation (APT, or SNPE-C) is also taken from the sbi package; as a sequential method, the simulation budget per round is computed by dividing the total budget by the number of rounds. For time series problems, a summary network is recommended for NPE/APT [Zeghal et al., 2022]. Accordingly, we modify and optimize the standard CNNEmbedding from the sbi package to convert the data into summary statistics with dimensions matching the parameters (4 for the LV test problem and 2 for the SIR task).
For the ConDiSim model, we optimized the hyperparameters for each of the ten benchmark problems from the sbibm library, as detailed in Table 3. Common hyperparameters were applied uniformly across all tasks to ensure training consistency and stability. We used the AdamW optimizer with a learning rate of $0.001$ and a weight decay of $0.0001$, along with a ReduceLROnPlateau scheduler to decrease the learning rate by a factor of $0.5$ if validation loss did not improve over five consecutive epochs. The noise schedule followed a cosine/quadratic scheme, with $\beta_{\text{min}}$ as $0.0001$ and $\beta_{\text{max}}$ as $0.02$, to regulate progressive noise addition during diffusion. Data standardization was conducted using StandardScaler, and an $80/20$ split was used for training and validation. Early stopping with a patience of $20$ epochs was applied to prevent overfitting. SiLU (Sigmoid Linear Unit) activations were used across all layers to enhance pattern capture and improve stability. This consistent training strategy enabled a systematic comparison of model performance across various batch sizes, diffusion steps, and hidden unit configurations.
Table 3: Hyperparameters for ConDiSim Model
@@
Benchmark Batch Size Diffusion Steps Hidden Units SLCP 32 200 Linear, (512, 512, 512) SLCP with Distractors 32 1000 Linear, (512, 512, 512) Two Moons 32 50 Linear, (256, 256, 256) Gaussian Mixture 32 50 Linear, (256, 256, 256) Gaussian Linear Uniform 32 50 Linear, (256, 256, 256) SIR 32 50 Linear, (256, 256, 256) Bernoulli GLM 128 500 Linear, (256, 256, 256) Bernoulli GLM Raw 128 500 Linear, (512, 512, 512) Lotka-Volterra 128 200 Conv1D (32, 64, 128), Linear, (128, 128) HH Model 32 300 Linear, (256, 256, 256) Genetic Oscillator 64 400 Linear, (256, 256, 256, 256)
Appendix C Additional Results
We present plots of estimated posterior distributions for the remaining benchmark test problems below:
C.1 Posterior Evolution for Two Moons
@@
(a) Hodgkin–Huxley: Approximated Posterior
(b) Posterior Predictive Voltage Traces
Figure 9: Simformer:Posterior Distributions for Hodgkin–Huxley model
C.2 Performance of Simformer on HH Model
Unlike the standard SBI benchmarks, where true posteriors are known for a known set of true parameters, the Hodgkin-Huxley (HH) model lacks an analytical posterior, preventing direct comparison via metrics like C2ST or MMD. For fair evaluation, we first benchmark CondiSim on 10 standard SBI tasks (refer Section 4.1) before demonstrating its real-world applicability to the HH model. Given a fixed parameter configuration $\theta^{*}=(1.437,72.177,24.209,0.154,66.87,-81.435,-67.606)$, we infer posteriors using both CondiSim and Simformer (trained with 30k simulations). Posterior samples and posterior predictive checks for Simformer (Figure 9) and ConDiSim (Figure 4) confirm that both methods produce plausible voltage traces, with ConDiSim yielding tighter posterior support on several parameters.
Appendix D Posterior Plots for SBI benchmarking problems
In this section, we summarize the unique challenges posed by remaining SBI benchmark problem and highlight the performance of our model. Detailed posterior distributions are provided in the corresponding figures for visual reference.
@@ Appendix D Posterior Plots for SBI benchmarking problems
For Gaussian Linear and Gaussian Linear Uniform, the challenge is accurately modeling the posterior under both conjugate and non-conjugate settings. The Gaussian Linear Uniform variant complicates posterior inference with a uniform prior. Figure 11(a) and Figure 11(b) shows that our model effectively handles these differences.
In Bernoulli GLM and Bernoulli GLM Raw, the challenge involves estimating posteriors from Bernoulli observations with Gaussian priors. The "Raw" variant introduces complexity by using raw data instead of sufficient statistics. Figures 12(a) and 12(b) indicate accurate posterior modeling for both cases.
The SIR model aims to infer contact and recovery rates from noisy infection data over time. Figure 10(b) illustrates that our model robustly estimates the parameters despite the high variability in data.
@@
(a) Gaussian Linear Model
(b) Gaussian Linear Uniform Model
Figure 11: Posterior Distributions for Gaussian Linear and Gaussian Linear Uniform Models: Generated vs. Reference Posterior
(a) Bernoulli GLM
(b) Bernoulli GLM Raw
Figure 12: Posterior Distributions for Bernoulli GLM and Bernoulli GLM Raw: Generated vs. Reference Posterior
Figure 13: Lotka–Volterra Posterior: Visualization of the posterior for the Lotka–Volterra model. The true posterior mean is indicated by a vertical line (due to low variance).
Appendix E Simulation Based Calibration
Simulation-Based Calibration (SBC) is an essential metric for generative models in Simulation-Based Inference (SBI), as it verifies alignment between inferred posterior distributions and true parameter values, assisting in detecting potential biases in model-generated samples. For SBC, we generate $M=500$ prior samples, simulate corresponding observations, and draw $L=250$ posterior samples for each observation using the learned ConDiSim model. Calibration is assessed through rank statistics of true parameters within their posterior distributions, with empirical cumulative distribution functions (ECDFs) providing a robust, bin-free visualization of these ranks. Following [Säilynoja et al., 2022], ECDF-based SBC transforms posterior draws into fractional rank statistics, avoiding arbitrary binning and facilitating interpretation. A uniform ECDF indicates well-calibrated posteriors, while deviations suggest miscalibration: curves above the diagonal indicate overconfidence, and those below indicate underconfidence. Using BayesFlow [Radev et al., 2023b], we generate ECDF plots with 90% confidence interval bands as calibration benchmarks. Solid blue lines represent rank ECDFs of posterior draws, with shaded grey areas denoting the 90% confidence intervals. The SBC results for SBI benchmarking problems are presented below:
(a) SLCP
(b) SLCP Distractors
(c) Gaussian Linear
(d) Gaussian Linear Uniform
(e) Bernoulli GLM
(f) Bernoulli GLM Raw
(g) SIR
(h) Gaussian Mixture
(i) Lotka Volterra
@@ Appendix E Simulation Based Calibration
Appendix F SBI Metrics
F.1 Limitations of the C2ST Metric
The Classifier Two-Sample Test (C2ST) evaluates the similarity between a generated posterior $P$ and the true posterior $Q$ by training a classifier to differentiate their samples. A score near 0.5 indicates that the two distributions are similar, while higher values indicate discrepancies. Although widely used in SBI literature for its straightforward interpretability—providing a clear score between 0.5 and 1—the metric can be overly sensitive, as a single poorly estimated dimension may distort the results. Furthermore, its reliability depends significantly on the design and training of the classifier.
F.2 MMD
To overcome these limitations, we also consider the Maximum Mean Discrepancy (MMD), which measures the distance between the mean embeddings of $P$ and $Q$ in a reproducing kernel Hilbert space (RKHS). MMD is defined as,
@@ F.2 MMD
where $\phi$ is a feature mapping into the RKHS $\mathcal{H}$. A low MMD value indicates that the generated posterior closely approximates the true posterior. Unlike C2ST, MMD provides a robust, holistic measure of similarity between distributions without being overly influenced by discrepancies in a single dimension. A low MMD value indicates that the distributions $P$ and $Q$ are similar, while a larger value suggests significant differences.
@@
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.1213 $\pm$0.0650 0.0582 $\pm$0.0122 0.1432 $\pm$0.1161 3.6381 $\pm$0.5428 4.8299 $\pm$0.6825 3.9965 $\pm$1.8510 2.9470 $\pm$0.4891 2.7502 $\pm$0.7268 2.1466 $\pm$0.5625 NPE 0.0254 $\pm$0.0025 0.0241 $\pm$0.0099 0.0125 $\pm$0.0045 1.0371 $\pm$0.1830 0.9379 $\pm$0.1001 0.8675 $\pm$0.1945 0.6611 $\pm$0.2236 1.0312 $\pm$0.1255 1.1903 $\pm$0.1007 APT 0.0073 $\pm$0.0030 0.0069 $\pm$0.0030 0.0037 $\pm$0.0012 0.7383 $\pm$0.0853 0.4944 $\pm$0.0663 0.5808 $\pm$0.0539 1.1003 $\pm$0.0712 1.1093 $\pm$0.0311 1.1291 $\pm$0.0481 Simformer 0.0005 $\pm$0.0003 0.0001 $\pm$0.0001 0.0003 $\pm$0.0004 0.0191 $\pm$0.0098 0.0156 $\pm$0.0087 0.0116 $\pm$0.0061 0.0018 $\pm$0.0028 0.0016 $\pm$0.0042 0.0016 $\pm$0.0085 FMPE 0.0133 $\pm$ 0.0197 0.0039 $\pm$ 0.0060 0.019 $\pm$ 0.0014 0.1076 $\pm$ 0.0283 0.0826 $\pm$ 0.0138 0.0746 $\pm$ 0.0247 0.2058 $\pm$ 0.0024 0.3618 $\pm$ 0.0048 0.1819 $\pm$ 0.0044 ConDiSim 0.0105 $\pm$0.0041 0.0104 $\pm$0.0048 0.0074 $\pm$0.0002 0.3157 $\pm$0.1771 0.1471 $\pm$0.0428 0.1981 $\pm$0.1348 0.7560 $\pm$0.1803 0.5754 $\pm$0.2392 0.6909 $\pm$0.1521
(b) SLCP Distractors, Bernoulli GLM Raw, and Lotka Volterra
@@
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 5.1993 $\pm$1.2320 7.2478 $\pm$1.2260 3.0631 $\pm$0.3244 3.4369 $\pm$0.8517 4.3479 $\pm$0.5433 3.5950 $\pm$0.8613 7.2684 $\pm$0.5733 5.5183 $\pm$1.7126 4.8858 $\pm$1.2146 NPE 1.2167 $\pm$0.2397 0.7834 $\pm$0.3453 0.6428 $\pm$0.1779 1.1227 $\pm$0.0987 1.1389 $\pm$0.1878 1.1448 $\pm$0.1287 2.6357 $\pm$0.3468 3.8635 $\pm$0.2770 4.3310 $\pm$0.0709 APT 1.8380 $\pm$0.4448 0.8278 $\pm$0.2373 0.4422 $\pm$0.1428 0.7233 $\pm$0.2131 0.6968 $\pm$0.0826 0.5437 $\pm$0.1333 3.1921 $\pm$0.8198 3.2722 $\pm$0.8169 4.4739 $\pm$0.9642 Simformer 0.0540 $\pm$0.0173 0.0541 $\pm$0.0161 0.0552 $\pm$0.0195 0.2056 $\pm$0.0486 0.0926 $\pm$0.0493 0.0731 $\pm$0.0238 0.9696 $\pm$0.4434 0.8342 $\pm$0.4742 0.9634 $\pm$0.2588 FMPE 0.1598 $\pm$ 0.0501 0.1403 $\pm$ 0.0471 0.0930 $\pm$ 0.0561 0.0859 $\pm$ 0.0167 0.0831 $\pm$ 0.0181 0.0781 $\pm$ 0.0102 1.0537 $\pm$ 0.6277 1.0858 $\pm$ 0.5639 1.1696 $\pm$ 0.4336 ConDiSim 1.0688 $\pm$0.3057 0.6939 $\pm$0.3204 0.4105 $\pm$0.3082 0.1448 $\pm$0.0368 0.1435 $\pm$0.0822 0.1658 $\pm$0.1407 3.4480 $\pm$0.8115 3.4678 $\pm$0.3850 3.7210 $\pm$0.5691
(c) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.2242 $\pm$0.0296 0.1149 $\pm$0.0549 0.1733 $\pm$0.1082 0.4999 $\pm$0.1595 0.6476 $\pm$0.3050 0.7182 $\pm$0.2839 0.9044 $\pm$0.2353 0.7712 $\pm$0.1483 1.1197 $\pm$0.4795 0.7197 $\pm$0.5598 0.3375 $\pm$0.1473 0.7948 $\pm$0.5992 NPE 0.0604 $\pm$0.0260 0.0418 $\pm$0.0211 0.0554 $\pm$0.0347 0.0105 $\pm$0.0036 0.0074 $\pm$0.0018 0.0082 $\pm$0.0011 0.0258 $\pm$0.0098 0.0117 $\pm$0.0041 0.0121 $\pm$0.0051 0.4021 $\pm$0.0551 0.1632 $\pm$0.0684 0.2333 $\pm$0.0935 APT 0.0137 $\pm$0.0031 0.0108 $\pm$0.0088 0.0110 $\pm$0.0034 0.0075 $\pm$0.0013 0.0070 $\pm$0.0018 0.0080 $\pm$0.0008 0.0166 $\pm$0.0087 0.0138 $\pm$0.0053 0.0113 $\pm$0.0039 0.1405 $\pm$0.0347 0.0487 $\pm$0.0296 0.0287 $\pm$0.0166 Simformer 0.0009 $\pm$0.0007 0.0009 $\pm$0.0006 0.0007 $\pm$0.0010 0.0092 $\pm$0.0116 0.0037 $\pm$0.0035 0.0037 $\pm$0.0028 0.0014 $\pm$0.0006 0.0015 $\pm$0.0011 0.0008 $\pm$0.0003 0.0323 $\pm$0.0203 0.0025 $\pm$0.0034 0.0021 $\pm$0.0030 FMPE 0.1893 $\pm$ 0.0985 0.0208 $\pm$ 0.0078 0.0013 $\pm$ 0.0011 0.0511 $\pm$ 0.0053 $0$$\pm$ 0.0041 0.0497 $\pm$ 0.0034 0.1425 $\pm$ 0.0219 0.1261 $\pm$ 0.0206 0.1236 $\pm$ 0.0247 0.1573 $\pm$ 0.0543 0.1489 $\pm$ 0.0528 0.1489 $\pm$ 0.0528 ConDiSim 0.1604 $\pm$0.0599 0.0720 $\pm$0.0124 0.0974 $\pm$0.0505 0.0471 $\pm$0.0126 0.0387 $\pm$0.0129 0.0290 $\pm$0.0038 0.0335$\pm$0.0139 0.0273 $\pm$0.0079 0.0285 $\pm$0.0098 0.2760 $\pm$0.0493 0.1954 $\pm$0.1029 0.1572 $\pm$0.0726

### v2

@@ B.11 Hodgkin-Huxley Model
  • Fixed Parameters: The initial membrane voltage is $V_{0}=-65.0\,\text{mV}$, with gating variables initialized to their steady-state values at $V_{0}$. Simulations use a time resolution of $5000\,\text{steps}$ over $200\,\text{ms}$ simulation window.
B.12 Genetic Oscillator
The genetic oscillator model describes a biochemical reaction network involving $9$ species governed by $18$ reactions, parameterized by $15$ reaction rate constants. This network is based on a positive–negative feedback loop, where the activator protein $A$ upregulates transcription of both its gene and the repressor protein $R$. The repressor $R$ sequesters $A$ by forming a complex $C$, thereby inhibiting further activation [Vilar et al., 2002, Akesson et al., 2022]. The species in the network are:
@@ B.12 Genetic Oscillator
The network includes the following reactions:
$D_{A},A$ $\xrightarrow{\gamma_{A}}D_{A}^{*},$ $D_{A}^{*}$ $\xrightarrow{\theta_{A}}D_{A},$
$D_{A}^{*}$ $\xrightarrow{\theta_{A}}D_{A},\,A,$ $D_{A}^{*}$ $\xrightarrow{\alpha_{A}^{*}}D_{A}^{*},\,M_{A},$
$D_{A}$ $\xrightarrow{\alpha_{A}}D_{A},\,M_{A},$
$D_{R},A$ $\xrightarrow{\gamma_{R}}D_{R}^{*},$ $D_{R}^{*}$ $\xrightarrow{\theta_{R}}D_{R},$
$D_{R}^{*}$ $\xrightarrow{\theta_{A}}D_{R}^{*},\,A,$ $D_{R}^{*}$ $\xrightarrow{\alpha_{R}^{*}}D_{R}^{*},\,M_{R},$
$D_{R}$ $\xrightarrow{\alpha_{R}}D_{R},\,M_{R},$
@@
[…]
@@
  • Simulation Time: The system is simulated over $200$ time steps uniformly distributed across $200\,\text{ms}$.
Appendix C Hyperparameters
In this section, we describe the hyperparameters and simulation environments used across all methods. Training data for the SBI benchmark problems were generated using sbibm [Lueckmann et al., 2021], except for the Lotka–Volterra (LV) and SIR models, which were simulated with BayesFlow [Radev et al., 2023b]. For SimFormer, we adopt the hyperparameters reported in Gloeckler et al. [2024] for both the general SBI setting and the Hodgkin–Huxley (HH) model. For the Genetic Oscillator model, we follow the simulation and preprocessing setup of Akesson et al. [2022]. Specifically, SimFormer is trained on a simulation budget of 30,000 using an 8-layer Transformer with 8 attention heads, 64-dimensional head projections, a 512-unit MLP, and dropout rate 0.1; it is optimized with Adam (learning rate $10^{-3}$) for 100,000 iterations with batch size 128, and posterior samples are generated via Euler–Maruyama integration over 1,000 diffusion steps. For GATSBI, we start from the task-specific hyperparameters in Ramesh et al. [2022] and adjust them as needed based on problem structure and computational constraints. All experiments were run on a system with dual 16-core Intel Xeon Gold 6226R CPUs, 576 GB RAM, and an NVIDIA Tesla T4 GPU (16 GB).
@@ Appendix C Hyperparameters
For NPE, we adopt the setup of Lueckmann et al. [2021], who evaluate NPE across ten sbibm benchmark tasks. Appendix H of their work shows that Neural Spline Flows (NSF) consistently outperform Masked Autoregressive Flows (MAF), which were originally proposed as the density estimator for NPE. We therefore use the NSF implementation from the sbi package [Tejero-Cantero et al., 2020] in all experiments. We also use sbi’s implementation of Automatic Posterior Transformation (APT, or SNPE-C). As a sequential method, APT allocates the total simulation budget uniformly across rounds.
For ConDiSim, we tuned hyperparameters individually for each task in the sbibm benchmark suite [Lueckmann et al., 2021] while keeping the training pipeline fixed. We optimized with AdamW [Loshchilov and Hutter, 2019] using a YOLOX-inspired learning rate schedule: a quadratic warmup over the first $10\%$ of steps from $0.2\eta$ to the base learning rate $\eta$, followed by cosine annealing down to $10^{-6}$, and a constant phase at $10^{-6}$ for the final $5\%$ of steps; weight decay was set to $\lambda=10^{-4}$. We use scaled linear and scaled quadratic noise schedule [Ho et al., 2020] with $\beta_{\text{start}}=10^{-4}s$ and $\beta_{\text{end}}=0.02s$ with $s=1000/T$ to maintain consistent noise levels across varying diffusion steps $T$. Inputs were standardized to zero mean and unit variance, and a $70/30$ train–validation split was employed to handle the high variance of posterior estimation. Training stability was ensured via gradient clipping (max norm $5.0$) and SNR-based loss weighting [Hang et al., 2023] with $\gamma=5.0$, where timestep weights $w_{t}=\min(\mathrm{SNR}_{t},\gamma)/\mathrm{SNR}_{t}$ and $\mathrm{SNR}_{t}=\bar{\alpha}_{t}/(1-\bar{\alpha}_{t})$ emphasize learning at noisier timesteps. Early stopping (patience $30$ after epoch $50$) and SiLU activations throughout the network were also used. Task-specific hyperparameters are provided in Table 3.
Table 3: ConDiSim hyperparameters.
@@
Benchmark # Diffusion Blocks Hidden Dim Scheduler # Timesteps Batch Size Learning Rate SLCP 6 128 linear 200 32 1e-3 SLCP with Distractors 6 128 quadratic 1000 50 1e-3 Two Moons 4 64 cosine 160 32 1e-3 Gaussian Mixture 4 64 cosine 160 32 1e-3 Gaussian Linear 6 64 linear 100 50 2e-4 Gaussian Linear Uniform 6 64 cosine 200 50 2e-4 SIR 4 64 linear 100 32 1e-4 Bernoulli GLM 6 128 linear 200 32 1e-3 Bernoulli GLM Raw 6 128 linear 200 32 1e-3 Lotka-Volterra 6 128 cosine 200 50 2e-4 HH Model 6 128 quadratic 1000 50 1e-3 Genetic Oscillator 6 128 cosine 300 64 1e-4
Appendix D Additional Results
D.1 Posterior Evolution for Two Moons
@@
D.2 Comparison with Simformer on HH and Genetic Oscillator Models
Unlike standard SBI benchmarks with analytical posteriors, the Hodgkin–Huxley (HH) and Genetic Oscillator models lack closed-form solutions, preventing direct evaluation via C2ST or MMD. Having benchmarked ConDiSim on ten standard SBI tasks (Section 4.1), we now compare it against Simformer, the closest competing method, on these real-world problems.
Both methods are trained with 30k simulations. For the HH model with fixed parameters $\bm{\theta^{*}}=(1.437,72.177,24.209,0.154,66.87,-81.435,-67.606)$, posterior samples and predictive checks (Figures 84) show both methods recover plausible voltage traces, with ConDiSim yielding tighter posterior support for several parameters. For the genetic oscillator, posterior distributions for Simformer and ConDiSim are shown in Figures 9 and 5(a), respectively. Both produce reasonable posteriors, but ConDiSim achieves sharper concentration around true parameters for several rates (e.g., $\alpha_{r},\beta_{r},\gamma_{c},\theta_{a}$), while Simformer yields broader supports. This demonstrates ConDiSim’s more precise uncertainty quantification without sacrificing coverage, particularly effective for complex oscillatory dynamics.
(a) Hodgkin–Huxley: Approximated Posterior.
(b) Posterior Predictive Voltage Traces.
Figure 8: Simformer: Posterior Distributions for Hodgkin–Huxley model.
Figure 9: Simformer: Posterior Distributions for the Genetic Oscillator test problem.
Appendix E Posterior Plots for SBI benchmarking problems
In this section, we summarize the unique challenges posed by the remaining SBI benchmark problem and highlight the performance of our model. Detailed posterior distributions are provided in the corresponding figures for visual reference.
@@ Appendix E Posterior Plots for SBI benchmarking problems
For Gaussian Linear and Gaussian Linear Uniform, the challenge is accurately modeling the posterior under both conjugate and non-conjugate settings. The Gaussian Linear Uniform variant complicates posterior inference with a uniform prior. Figure 11(a) and Figure 11(b) show that our model effectively handles these differences.
In Bernoulli GLM and Bernoulli GLM Raw, the challenge involves estimating posteriors from Bernoulli observations with Gaussian priors. The “Raw” variant introduces complexity by using raw data instead of sufficient statistics. Figures 12(a) and 12(b) indicate accurate posterior modeling for both cases.
The SIR model aims to infer contact and recovery rates from noisy infection data over time. Figure 10(b) illustrates that our model robustly estimates the parameters despite the high variability in the data.
@@
(a) Gaussian Linear Model.
(b) Gaussian Linear Uniform Model.
Figure 11: Posterior Distributions for Gaussian Linear and Gaussian Linear Uniform Models: Generated vs. Reference Posterior.
(a) Bernoulli GLM.
(b) Bernoulli GLM Raw.
Figure 12: Posterior Distributions for Bernoulli GLM and Bernoulli GLM Raw: Generated vs. Reference Posterior.
Figure 13: Lotka Volterra Posterior: Visualization of the posterior for the Lotka–Volterra model. The true posterior mean is indicated by a vertical line (due to low variance).
Appendix F Simulation Based Calibration
Simulation-Based Calibration (SBC) is an important metric for generative models in SBI, as it verifies alignment between inferred posterior distributions and true parameter values, assisting in detecting potential biases in model-generated samples. For SBC, we generate $M=500$ prior samples, simulate corresponding observations, and draw $L=250$ posterior samples for each observation using the learned ConDiSim model. Calibration is assessed through rank statistics of true parameters within their posterior distributions, with empirical cumulative distribution functions (ECDFs) providing a robust, bin-free visualization of these ranks. Following [Säilynoja et al., 2022], ECDF-based SBC transforms posterior draws into fractional rank statistics, avoiding arbitrary binning and facilitating interpretation. A uniform ECDF indicates well-calibrated posteriors, while deviations suggest miscalibration: curves above the diagonal indicate overconfidence, and those below indicate underconfidence. Using BayesFlow [Radev et al., 2023b], we generate ECDF plots with 90% confidence interval bands as calibration benchmarks. Solid blue lines represent rank ECDFs of posterior draws, with shaded grey areas denoting the 90% confidence intervals. The SBC results for SBI benchmarking problems are presented below:
(a) SLCP.
(b) SLCP Distractors.
(c) Gaussian Linear.
(d) Gaussian Linear Uniform.
(e) Bernoulli GLM.
(f) Bernoulli GLM Raw.
(g) SIR.
(h) Gaussian Mixture.
(i) Lotka Volterra.
@@ Appendix F Simulation Based Calibration
Appendix G SBI Metrics
G.1 Limitations of the C2ST Metric
The Classifier Two-Sample Test (C2ST) evaluates the similarity between a generated posterior $P$ and the true posterior $Q$ by training a classifier to differentiate their samples. A score near 0.5 indicates that the two distributions are similar, while higher values indicate discrepancies. Although widely used in SBI literature for its straightforward interpretability providing a clear score between 0.5 and 1, the metric can be overly sensitive, as a single poorly estimated dimension may distort the results. Furthermore, its reliability depends significantly on the design and training of the classifier [Friedman, 2003, Lopez-Paz and Oquab, 2016]. However, C2ST remains the most widely used metric in SBI, as it is both interpretable and well established for evaluating benchmark problems. For completeness, we also report MMD results in the following section, although MMD is kernel dependent, less commonly used, and often difficult to interpret solely from numerical values.
G.2 MMD
The Maximum Mean Discrepancy (MMD) measures the distance between the mean embeddings of $P$ and $Q$ in a reproducing kernel Hilbert space (RKHS). MMD is defined as,
@@ G.2 MMD
where $\phi$ is a feature mapping into the RKHS $\mathcal{H}$. A low MMD value indicates that the generated posterior closely approximates the true posterior. MMD therefore provides a holistic measure of similarity between distributions without being overly influenced by discrepancies in a single dimension. A low MMD value indicates that the distributions $P$ and $Q$ are similar, while a larger value suggests significant differences.
@@
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 0.0254 $\pm$0.0025 0.0241 $\pm$0.0099 0.0125 $\pm$0.0045 1.0371 $\pm$0.1830 0.9379 $\pm$0.1001 0.8675 $\pm$0.1945 0.6611 $\pm$0.2236 1.0312 $\pm$0.1255 1.1903 $\pm$0.1007 APT 0.0073 $\pm$0.0030 0.0069 $\pm$0.0030 0.0037 $\pm$0.0012 0.7383 $\pm$0.0853 0.4944 $\pm$0.0663 0.5808 $\pm$0.0539 1.1003 $\pm$0.0712 1.1093 $\pm$0.0311 1.1291 $\pm$0.0481 GATSBI 0.1213 $\pm$0.0650 0.0582 $\pm$0.0122 0.1432 $\pm$0.1161 3.6381 $\pm$0.5428 4.8299 $\pm$0.6825 3.9965 $\pm$1.8510 2.9470 $\pm$0.4891 2.7502 $\pm$0.7268 2.1466 $\pm$0.5625 Simformer 0.0005 $\pm$0.0003 0.0001 $\pm$0.0001 0.0003 $\pm$0.0004 0.0191 $\pm$0.0098 0.0156 $\pm$0.0087 0.0116 $\pm$0.0061 0.0018 $\pm$0.0028 0.0016 $\pm$0.0042 0.0016 $\pm$0.0085 FMPE 0.0133 $\pm$ 0.0197 0.0039 $\pm$ 0.0060 0.019 $\pm$ 0.0014 0.1076 $\pm$ 0.0283 0.0826 $\pm$ 0.0138 0.0746 $\pm$ 0.0247 0.2058 $\pm$ 0.0024 0.3618 $\pm$ 0.0048 0.1819 $\pm$ 0.0044 ConDiSim 0.0172 $\pm$0.0226 0.0091 $\pm$0.0099 0.0052 $\pm$0.0018 0.0393 $\pm$0.0095 0.0525 $\pm$0.0547 0.0190 $\pm$0.0134 0.2413 $\pm$0.0681 0.3853 $\pm$0.1195 0.4860 $\pm$0.1198
(b) SLCP Distractors, Bernoulli GLM Raw, and Lotka Volterra.
@@
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 1.2167 $\pm$0.2397 0.7834 $\pm$0.3453 0.6428 $\pm$0.1779 1.1227 $\pm$0.0987 1.1389 $\pm$0.1878 1.1448 $\pm$0.1287 2.6357 $\pm$0.3468 3.8635 $\pm$0.2770 4.3310 $\pm$0.0709 APT 1.8380 $\pm$0.4448 0.8278 $\pm$0.2373 0.4422 $\pm$0.1428 0.7233 $\pm$0.2131 0.6968 $\pm$0.0826 0.5437 $\pm$0.1333 3.1921 $\pm$0.8198 3.2722 $\pm$0.8169 4.4739 $\pm$0.9642 GATSBI 5.1993 $\pm$1.2320 7.2478 $\pm$1.2260 3.0631 $\pm$0.3244 3.4369 $\pm$0.8517 4.3479 $\pm$0.5433 3.5950 $\pm$0.8613 7.2684 $\pm$0.5733 5.5183 $\pm$1.7126 4.8858 $\pm$1.2146 Simformer 0.0540 $\pm$0.0173 0.0541 $\pm$0.0161 0.0552 $\pm$0.0195 0.2056 $\pm$0.0486 0.0926 $\pm$0.0493 0.0731 $\pm$0.0238 0.9696 $\pm$0.4434 0.8342 $\pm$0.4742 0.9634 $\pm$0.2588 FMPE 0.1598 $\pm$ 0.0501 0.1403 $\pm$ 0.0471 0.0930 $\pm$ 0.0561 0.0859 $\pm$ 0.0167 0.0831 $\pm$ 0.0181 0.0781 $\pm$ 0.0102 1.0537 $\pm$ 0.6277 1.0858 $\pm$ 0.5639 1.1696 $\pm$ 0.4336 ConDiSim 0.1378 $\pm$0.0404 0.0580 $\pm$0.0205 0.0275 $\pm$0.0182 0.0968 $\pm$0.1012 0.0359 $\pm$0.0212 0.0305 $\pm$0.0279 2.3226 $\pm$0.1486 2.1965 $\pm$0.3856 2.1180 $\pm$0.2584
(c) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP.
[…]